# On the limitation of evaluating machine unlearning using only a single training seed

### v2

@@ Abstract
Machine unlearning (MU) aims to remove the influence of certain data points from a trained model without costly retraining. Most practical MU algorithms are only approximate and their performance can only be assessed empirically. Care must therefore be taken to make empirical comparisons as representative as possible. A common practice is to run the MU algorithm multiple times independently starting from the same trained model. In this work, we demonstrate that this practice can give highly non-representative results because – even for the same architecture and same dataset – some MU methods can be highly sensitive to the choice of random number seed used for model training. We therefore recommend that empirical comparisons of MU algorithms should also reflect the variability across different model training seeds.
@@
Figure 1: Retain-set accuracy (Panel 1(a)) and forget-set accuracy (Panel 1(b)), when using one training seed (combined with eleven unlearning seeds) versus eleven training seeds (each paired with one unlearning seed). Black lines represent the 25, 50 and 75 % quantiles of the accuracy of Retrain (the gold standard) across eleven seeds; whiskers extend to the full data range for clarity. This figure shows that the performance of some unlearning methods (e.g., SSD and LFSSD) is highly sensitive to choice of the training seed. Thus, empirical evaluation of unlearning methods should be based on multiple training seeds.
@@ 1 Introduction
Machine unlearning. The targeted removal of the influence of certain training data on the parameters of machine-learning models is often needed, e.g., for regulatory compliance (e.g., GDPR [10]), elimination of biased or corrupted samples, or mitigation of data poisoning attacks. These concerns extend to information retrieval systems [7, 8, 13]. However, the exact (‘gold-standard’) solution to the machine-unlearning (MU) problem, which involves retraining the model from scratch without the to-be-forgotten data (called Retrain), is typically infeasible. Instead, a number of approximate MU methods have been developed to enable models to forget specific data whilst maintaining overall performance, e.g., Random-Labels [5], UNSIR [12], Bad-Teacher [2], SSD [3] or LFSSD [4].
Evaluating unlearning methods. MU algorithms are often difficult to analyse mathematically. Their evaluation must then be done empirically. This calls for empirical experiments be designed carefully so as to provide confidence that their findings generalise beyond the specific setting of the study (e.g., architecture or dataset). For instance, common practice is to run the unlearning algorithm multiple times independently – each using a different random-number generator (RNG) seed, henceforth called an unlearning seed. One then assesses the empirical distribution (across unlearning seeds) of unlearning metrics such as accuracy on the data to be forgotten (forget-set accuracy) or accuracy on the remaining data (retain-set accuracy).
Contributions. While comparisons across multiple unlearning seeds are common, most works only consider a single training seed, i.e., they take the same single trained model as the starting point of unlearning. However, as we point out in this work, this practice can give highly non-representative results (see Figure 1). That is, even for the same architecture and dataset, some MU methods can be highly sensitive to the choice of training seed. This issue seems to particularly affect the deterministic MU methods: SSD and LFSSD. For a fixed trained model, such deterministic methods always return exactly the same unlearned model – and hence exactly the same values of any unlearning metric – no matter what unlearning seed is used. However, Figure 1 illustrates that unlearning metrics for SSD and LFSSD are actually highly variable across different training seeds. Specifically, our contributions are as follows:
@@
    We illustrate that single-training-seed based evaluation of unlearning methods produces non-representative results, especially for SSD and LFSSD.
@@
    We explain the counter-intuitive fact that an unlearning metric can exhibit a smaller variability across multiple training seeds than for a single, fixed training seed.
@@
    We recommend evaluating unlearning methods across multiple training seeds.
@@ 2 Methods
    • $J\geq 1$ unlearning seeds (e.g., J=3J=3[5], $J=5$ [2], $J=10$ [4, 3]).
@@
    Our recommendation is to evaluate unlearning across
@@ 3 Experiments
In this section, we demonstrate that the differences between the common-practice evaluation approach (A) and our recommended approach (B) shown in Figure 1 extends to other classes and datasets.
As datasets, we chose CIFAR-100 full-class [9], CIFAR-20 full-class, and CIFAR-20 sub-class [2]. For CIFAR-20 sub-class, the 100 classes in CIFAR-100 were grouped into 20 superclasses. We used a ResNet-18 [6] architecture, randomly initialised and trained from scratch on each dataset. Experiments were run on an NVIDIA A100-SMX4-80GB GPU. Code for reproducing the experiments – based on https://github.com/if-loops/selective-synaptic-dampening with minimal modifications – is available at https://github.com/jtlan90/evaluating-machine-unlearning-using-only-a-single-training-seed/.
@@ 4 Conclusion
Summary. We have demonstrated that the performance of machine unlearning (MU) methods – as measured by the accuracy on the retain and forget sets – can be highly sensitive to the initial model weights as determined by training seed. Thus, using only a single training seed for evaluating MU methods (which is common practice in the literature) can yield misleading conclusions about their efficacy or robustness. Results for another MU metric: Membership inference attack (MIA) [1, 11], were qualitatively similar and were thus omitted in the interest of brevity.
Recommendations. Based on our findings, MU evaluations should use multiple training seeds (>1>1) to ensure results are robust.
Limitations and extensions. Our work illustrates the potential pitfall of evaluating machine unlearning using only a single training seed. However, we note that for each dataset and model, our experiments compare only one instance of the ‘common-practice’ MU evaluation approach (i.e., a single training seed but multiple unlearning seeds) with only one instance of our recommended evaluation approach (i.e., multiple training seeds). A more comprehensive analysis (which would require multiple instances of each approach) is beyond the scope of this work. We leave this to future work.

### v3

@@ Abstract
Machine unlearning (MU) aims to remove the influence of certain data points from a trained model without costly retraining. Most practical MU algorithms are only approximate and their performance can only be assessed empirically. Care must therefore be taken to make empirical comparisons as representative as possible. A common practice is to run the MU algorithm multiple times independently starting from the same trained model. In this work, we demonstrate that this practice can give highly non-representative results because – even for the same architecture and same dataset – some MU methods can be highly sensitive to the choice of random number seed used for model training. We illustrate that this is particularly relevant for MU methods that are deterministic, i.e., which always produce the same result when started from the same trained model. We therefore recommend that empirical comparisons of MU algorithms should also reflect the variability across different model training seeds.
@@
Figure 1: Retain-set accuracy (Panel 1(a)) and forget-set accuracy (Panel 1(b)), when using one training seed (combined with eleven unlearning seeds) versus eleven training seeds (each paired with one unlearning seed). Black lines represent the 25, 50 and 75 % quantiles of the accuracy of Retrain (the gold standard) across eleven seeds; whiskers extend to the full data range for clarity. This figure shows that the performance of the deterministic unlearning methods: SSD and LFSSD, is highly sensitive to choice of the training seed. Thus, empirical evaluation of (especially such deterministic) unlearning methods should be based on multiple training seeds.
@@ 1 Introduction
Machine unlearning. The targeted removal of the influence of certain training data on the parameters of machine-learning models [Machine_Unlearning_Cao] is often needed, e.g., for regulatory compliance (e.g., GDPR [mantelero2013eu]), elimination of biased or corrupted samples, or mitigation of data poisoning attacks. These concerns extend to information retrieval systems [hou2024neural, hou2024neural_corrective, zhang2022machine]. However, the exact (‘gold-standard’) solution to the machine-unlearning (MU) problem, which involves retraining the model from scratch without the to-be-forgotten data (called Retrain), is typically infeasible. Instead, a number of approximate MU methods have been developed to enable models to forget specific data whilst maintaining overall performance, e.g., Random Labels [GolatkarEternalsunshineoftheSpotless202], UNSIR [Tarun_2024_UNSIR], Bad Teacher [chundawat2023badteachinginduceforgetting], SSD [Foster_Schoepf_Brintrup_2024] or LFSSD [foster2024loss].
Evaluating unlearning methods. MU algorithms are often difficult to analyse mathematically. Their evaluation must then be done empirically. This calls for empirical experiments to be designed carefully so as to provide confidence that their findings generalise beyond the specific setting of the study (e.g., architecture or dataset).
Common practice [chundawat2023badteachinginduceforgetting, Chundawat_2023, graves2021amnesiac, GolatkarEternalsunshineoftheSpotless202, Tarun_2024_UNSIR, Foster_Schoepf_Brintrup_2024, foster2024loss, ForgettingOutsidetheboxgolatkar, mixedprivforgettinggolatkar, bourtoule2021machine, unrollingsgdanvith, deepregressionunlearningtarun23a, lee2025unlearningcomparatorvisualanalytics] is to run stochastic unlearning algorithms (e.g., UNSIR, Random Labels or Bad Teacher) multiple times independently – each using a different random-number generator (RNG) seed, henceforth called an unlearning seed, to account for their inherent variability. One then assesses the empirical distribution (across unlearning seeds) of unlearning metrics such as accuracy on the data to be forgotten (forget-set accuracy) or accuracy on the remaining data (retain-set accuracy).
In contrast, deterministic MU algorithms, such as SSD and LFSSD, are only run once [Foster_Schoepf_Brintrup_2024, foster2024loss] because they always return exactly the same unlearned model – and hence exactly the same values of any unlearning metric – no matter what unlearning seed is used.
Contributions. While comparisons across multiple unlearning seeds are common, many works use a single training seed, i.e., a single trained model as the starting point of unlearning. In this work, we demonstrate that such a strategy can lead to highly non-representative results because – even for a given architecture and data set – MU methods can be highly sensitive to the choice of training seed. Our work thus confirms similar conclusions recently drawn in [triantafillou2024makingprogressunlearningfindings, cadet2025deep].
Notably, our work differs from [cadet2025deep, triantafillou2024makingprogressunlearningfindings] by demonstrating that this issue particularly affects the deterministic MU methods: SSD and LFSSD. As illustrated in Figure 1, unlearning metrics for SSD and LFSSD – although constant across unlearning seeds for a given training seed – exhibit disproportionately high variability across different training seeds.
Specifically, our contributions are as follows:
@@
    We illustrate that single-training-seed based evaluation of unlearning methods produces non-representative results, especially for the deterministic methods: SSD and LFSSD.
@@
    In particular, we show that using multiple unlearning seeds cannot compensate for the lack of multiple training seeds.
@@
    We explain the counter-intuitive fact that an unlearning metric can exhibit a smaller variability across multiple training seeds than for a single, fixed training seed (when the latter is combined with multiple unlearning seeds).
@@ 2 Methods
    • $J\geq 1$ unlearning seeds (e.g., J=3J=3[GolatkarEternalsunshineoftheSpotless202], $J=5$ [chundawat2023badteachinginduceforgetting], $J=10$ [foster2024loss, Foster_Schoepf_Brintrup_2024]).
@@
    Our recommendation reflecting that of [cadet2025deep] is to evaluate unlearning across
@@ 3 Experiments
In this section, we demonstrate that the differences between the common-practice evaluation approach (A) and our recommended approach (B) shown in Figure 1 extend to other classes and datasets.
As datasets, we chose CIFAR-100 full-class [krizhevsky2009learning], CIFAR-20 full-class, and CIFAR-20 sub-class [chundawat2023badteachinginduceforgetting]. For CIFAR-20 sub-class, the 100 classes in CIFAR-100 were grouped into 20 superclasses. We used a ResNet-18 [resnet] architecture, randomly initialised and trained from scratch on each dataset. Experiments were run on an NVIDIA A100-SMX4-80GB GPU. Code for reproducing the experiments – based on https://github.com/if-loops/selective-synaptic-dampening with minimal modifications – is available (anonymised for review) at https://anonymous.4open.science/r/evaluating-machine-unlearning-using-only-a-single-training-seed-88CC.
@@ 4 Conclusion
Summary. We have confirmed a recent finding from [cadet2025deep] that the performance of machine unlearning (MU) methods can be highly sensitive to the initial model weights as determined by training seed. Thus, using only a single training seed for evaluating MU methods (a common practice in the literature) can yield misleading conclusions about their efficacy or robustness. Complementary to [cadet2025deep], our work shows that: (a) this issue can be of particular concern when evaluating MU methods that are deterministic; (b) using multiple unlearning seeds cannot compensate for the lack of multiple training seeds.
Empirical results shown in this work are for the MU metrics: accuracy on the retain and forget sets; results for another MU metric: Membership inference attack (MIA) [MIAunlearningprivacychenetal, MIAshokrietal], were qualitatively similar and were thus omitted in the interest of brevity.
Recommendations. Based on our work we echo the recommendations of [cadet2025deep] that MU evaluations should use multiple training seeds (>1>1) to ensure results are robust.
Limitations and extensions. Our work specifically illustrates the potential pitfall of evaluating machine unlearning using only a single training seed. However, we note that for each dataset and model, our experiments compare only one instance of the ‘common-practice’ MU evaluation approach (i.e., a single training seed but multiple unlearning seeds) with only one instance of our recommended evaluation approach (i.e., multiple training seeds). A more comprehensive analysis (which would require multiple instances of each approach) is beyond the scope of this work. We leave this to future work.